# Shared Spatial Memory Through Predictive Coding

### v2

@@ Abstract
Constructing a consistent shared spatial memory is a critical challenge in multi-agent systems, where partial observability and limited bandwidth often lead to catastrophic failures in coordination. We introduce a multi-agent predictive coding framework that formulates coordination as the minimization of mutual uncertainty among agents. Through an information bottleneck objective, this framework prompts agents to learn not only who and what to communicate but also when. At the foundation of this framework lies a grid-cell-like metric as internal spatial coding for self-localization, emerging spontaneously from self-supervised motion prediction. Building upon this internal spatial code, agents gradually develop a bandwidth-efficient communication mechanism and specialized neural populations that encode partners’ locations—an artificial analogue of hippocampal social place cells (SPCs). These social representations are further utilized by a hierarchical reinforcement learning policy that actively explores to reduce joint uncertainty. On the Memory-Maze benchmark, our approach shows exceptional resilience to bandwidth constraints: success degrades gracefully from 73.5% to 64.4% as bandwidth shrinks from 128 to 4 bits/step, whereas a full-broadcast baseline collapses from 67.6% to 28.6%. Our findings establish a theoretically principled and biologically plausible basis for how complex social representations emerge from a unified predictive drive, leading to collective intelligence.
@@ Introduction
Collective intelligence enables groups to perform complex tasks exceeding individual capabilities [1, 2, 3, 4, 5, 6]. From ant colonies to wolf packs, forming and acting upon shared environmental understanding is paramount [7, 8, 9]. In spatial navigation, animals such as bats or rats coordinate movements by aligning internal spatial representations [10, 11, 12]. This alignment emerges through sparse exchange of high-level cues (sonar chirps, ultrasonic squeaks, visual signals), suggesting the existence of shared cognitive maps: distributed neural representations of environments including resources, hazards, and conspecifics [13, 14, 15, 16, 17].
Neural algorithms underlying shared representations extend innate mechanisms supporting individual cognition. In mammals, the hippocampal-entorhinal system provides the substrate for cognitive maps, with place cells encoding locations and grid cells providing a metric scaffold [18, 19, 20, 21, 22]. “Social place cells”—neurons firing when a partner occupies particular locations—offer compelling evidence for neural substrates integrating self with peers [23, 24, 25]. These findings indicate that the brain possesses sophisticated machinery not only for building a model of its own world but also for representing the world of others, a prerequisite for any meaningful social coordination. However, a critical distinction exists between representing others within single brains versus coordinating across brains through severely bandwidth-constrained communication. This raises a fundamental question: What computational principles enable agents to develop socially-aware representations and to coordinate these representations across distributed systems under limited communication?
Artificial multi-agent systems face the same challenge: the “communication bottleneck” [26, 27, 28, 29]. While promising for exploration, search, and rescue, collective performance degrades catastrophically when bandwidth is limited [30, 31, 32]. The core issue is inefficient management of statistical redundancy among agents. Concurrent exploration yields highly correlated sensory observations; naive strategies (broadcasting raw or compressed data) waste bandwidth transmitting redundant information partners could infer [33, 34, 35, 36]. How can decentralized systems learn a sparse yet optimally informative communication mechanism, transmitting only essential information to resolve partner uncertainty?
We demonstrate that shared spatial memory emerges from minimizing mutual predictive uncertainty between agents. Our multi-agent predictive coding framework implements this through: Level 1—grid-cell-like spatial metrics for self-localization; Level 2—bandwidth-efficient communication and emergent social place cells; Level 3—hierarchical policies reducing collective uncertainty through coordinated exploration.
@@
Unifying predictive coding and information bottleneck theory [10, 44], our work provides a theoretically grounded, biologically plausible basis for shared spatial memory. Our main contributions are as follows: ❶ A multi-agent framework enabling efficient, semantically rich communication from first principles. ❷ A computational model for social place cell emergence as necessary components of social predictive coding. ❸ Validation demonstrating state-of-the-art performance, scalability, and bandwidth insensitivity. This work enables collaborative systems coordinating with biological efficiency, grounded in computational principles underlying collective intelligence.
@@ Results
Having established that individual agents can build robust predictive models of their environment through grid-cell-like scaffolds, we next investigate how the framework extends this predictive principle to the social domain. A central challenge in multi-agent coordination is learning efficient communication under bandwidth constraints. While existing approaches rely on handcrafted communication protocols [26, 27] or learn to broadcast raw sensory data [30, 31], these methods struggle with bandwidth limitations and fail to discover semantically meaningful communication strategies. Unlike frameworks exploring asymmetric, teacher–student knowledge transfer [46], our setting addresses symmetric, peer-to-peer coordination. We compare our HRL-ICM framework against three baselines: No Communication, where agents search independently; Periodic Broadcast, where agents communicate at fixed intervals under bandwidth constraints; and Full Broadcast, where agents exchange information without bandwidth limitations. Framing communication through the information bottleneck principle, we demonstrate that agents can learn to cooperate by developing a communication mechanism guided by a singular objective: transmit only information that maximally reduces a partner’s future uncertainty. This simple requirement compels agents to collaboratively discover an emergent communication mechanism that is not only sparse and bandwidth-efficient, but also intelligent and semantically structured.
@@
In the previous section, we demonstrated that agents learn efficient communication through predictive coding. However, a more fundamental question remains: What internal representations must agents develop to effectively model and predict the states of their partners? While prior work has identified place cells encoding self-location in artificial agents [45, cueva2018emergence], the neural substrates for representing social partners in spatial contexts remain largely unexplored. Here, we show that a central challenge in collective intelligence—forming and maintaining representations of others—can be addressed through predictive learning.
@@ Methods
Z.F., Y.G. and J.W. conceived the predictive coding framework. Z.F., Y.G., Y.Z., and H.A. performed all experiments and data analysis. Y.G. and H.A. contributed to the conception of the HRL-ICM framework and data collection. J.W. and Y.F. supervised the entire project. Y.W. and Y.F. wrote and reviewed the manuscript. All authors discussed the results and provided critical feedback on the manuscript. J.W. and Y.F. provided funding.
@@
Zhengru Fang1,⋆, Yu Guo1,⋆, Jingjing Wang2,⋆, Yuang Zhang3, Haonan An1, Yinhai Wang3, Yuguang Fang1
1City University of Hong Kong, 2Beihang University, 3University of Washington, Equal contribution

### v4

@@ Abstract
The mammalian hippocampus harbors specialized neural populations, often termed social place cells, that encode the locations of conspecifics and support collective navigation. Although such representations are believed to arise from the same predictive machinery supporting individual spatial cognition, a computational principle that can reproduce this capacity in artificial multi-agent systems remains unclear. Here, we introduce a multi-agent predictive coding framework that casts coordination as the minimization of mutual predictive uncertainty, namely each agent reduces uncertainty about other agents’ future states under limited communication. This objective induces an information bottleneck that adaptively determines who communicates when and what to transmit, while self-supervised motion prediction promotes the emergence of grid-cell-like spatial codes. Notably, the same learning dynamics give rise to specialized neural units that encode partners’ locations, serving as artificial analogues of hippocampal social place cells (SPCs). In-silico lesion experiments further show that these social representations are causally required for coordination. On the Memory-Maze benchmark, our approach degrades gracefully from 73.5% to 64.4% success as bandwidth shrinks from 128 to 4 bits/step, whereas a full-broadcast baseline collapses from 67.6% to 28.6%. Together, these results identify predictive uncertainty minimization as a biologically grounded blueprint for bandwidth-efficient multi-agent coordination.
@@ Introduction
Collective intelligence enables groups to solve complex tasks exceeding individual capabilities [1, 2, 3, 4, 5, 6]. From ant colonies to wolf packs, effective coordination depends on forming and acting on shared environmental understanding [7, 8, 9]. Specifically, in spatial navigation, animals align internal representations through sparse exchange of high-level cues, suggesting the presence of shared cognitive maps: distributed representations that encode resources, hazards, and conspecifics [10, 11, 12, 13, 14, 15].
Neural algorithms underlying shared representations extend innate mechanisms supporting individual cognition. In mammals, the hippocampal-entorhinal system provides the substrate for cognitive maps, with place cells encoding locations and grid cells providing a metric scaffold [16, 17, 18, 19, 20]. “Social place cells”—neurons firing when a partner occupies particular locations—offer compelling evidence for neural substrates integrating self with peers [21, 22, 23]. These findings indicate that the brain possesses sophisticated machinery not only for building a model of its own world but also for representing the world of others, a prerequisite for any meaningful social coordination. However, a critical distinction exists between representing others within single brains versus coordinating across brains through severely bandwidth-constrained communication. This raises a fundamental question: What computational principle, grounded in how biological brains organize social cognition, would be sufficient for artificial agents to reproduce this capacity under severely bandwidth-constrained communication?
Existing multi-agent communication methods offer only partial answers because none treats communication as sending what a partner will need next [24]. Learned recipient-selection approaches [25, 26, 27] determine who receives a message, but do not optimize message content to reduce the receiver’s future uncertainty. Bandwidth-aware scheduling methods [28, 29, 30] determine when to transmit based on local triggers or heuristics, without considering the receiver’s predictive needs. Information-bottleneck approaches [31, 32, 33] compress messages to remove redundancy, but optimize for immediate relevance rather than downstream predictive utility, and performance degrades sharply under tight bandwidth constraints [34]. What all three families lack is an explicit model of the partner: without predicting what a partner will observe next, an agent cannot identify which information would most help that partner, which is precisely the strategy that makes biological coordination efficient.
We demonstrate that the optimal shared spatial memory emerges from minimizing mutual predictive uncertainty between agents. Our multi-agent predictive coding framework implements this through: Level 1, grid-cell-like spatial metrics for self-localization; Level 2, bandwidth-efficient communication and emergent social place cells; Level 3, hierarchical policies reducing collective uncertainty through coordinated exploration.
@@
Unifying predictive coding and information bottleneck theory [10, 42], our work provides a theoretically grounded, biologically plausible basis for shared spatial memory. Our main contributions are as follows: ❶ A computational model showing that social place cell-like representations emerge as a necessary consequence of a social predictive coding objective, offering a mechanistic account of their biological origins. ❷ A multi-agent framework enabling efficient, semantically rich communication from first principles. ❸ Validation demonstrating state-of-the-art performance, scalability, and bandwidth robustness in embodied cooperative navigation. Together, these results establish predictive uncertainty minimization as a bridge between individual spatial cognition and collective intelligence.
@@ Results
Having established that individual agents can build robust predictive models of their environment through grid-cell-like scaffolds, we next investigate how the framework extends this predictive principle to the social domain. A central challenge in multi-agent coordination is learning efficient communication under bandwidth constraints. Unlike frameworks exploring asymmetric, teacher–student knowledge transfer [44], our setting addresses symmetric, peer-to-peer coordination. We compare our HRL-ICM framework against three baselines: No Communication, where agents search independently; Periodic Broadcast, where agents communicate at fixed intervals under bandwidth constraints; and Full Broadcast, where agents exchange information without bandwidth limitations. Framing communication through the information bottleneck principle, we demonstrate that agents can learn to cooperate by developing a communication mechanism guided by a singular objective: transmit only information that maximally reduces a partner’s future uncertainty. This simple requirement compels agents to collaboratively discover an emergent communication mechanism that is not only sparse and bandwidth-efficient, but also intelligent and semantically structured.
@@
In the previous section, we demonstrated that agents learn efficient communication through predictive coding. However, a more fundamental question remains: What internal representations must agents develop to effectively model and predict the states of their partners? While prior work has identified place cells encoding self-location in artificial agents [43], the neural substrates for representing social partners in spatial contexts remain largely unexplored. Here, we show that a central challenge in collective intelligence—forming and maintaining representations of others—can be addressed through predictive learning.
@@ Methods
Z.F. and Y.G. conceived the predictive coding framework. Z.F., Y.G., Y.Z., and H.A. performed all experiments and data analysis. Y.G. and H.A. contributed to the conception of the HRL-ICM framework and data collection. W.D. and Y.F. supervised the entire project. W.D. and Y.F. wrote and reviewed the manuscript. All authors discussed the results and provided critical feedback on the manuscript. Y.F. provided funding.
@@
Zhengru Fang1,4,⋆, Yu Guo1,⋆, Jingjing Wang2,⋆, Yuang Zhang3, Haonan An1,
Yinhai Wang3, Wenbo Ding4, Yuguang Fang1
1City University of Hong Kong, 2Beihang University,
3University of Washington, 4Tsinghua University, Equal contribution